A man 1.5 m tall is on the top of a building .He observes a car on the road at an angle of 75°.if the building is 30 m high ,how far is the car from the building?

Answer by mananth(16946)


Total height of observation 30+1.5 = 31.5 m

tan 75 = 31.5/x
x= 31.5/ tan 75
x = 8.44 m